a tea shop owner wishes to blend P2.50/g tea with P3.25/g to produce a blend selling for P3/g.how much of each should be used to produce 7.5 kg of the blend

Answer by mananth(16946)

Price ---------------- quantity
Tea type I P2.50 ---------------- x gm
Tea type II P3.25 ------ 7500 - x gm
Mixture P3.00 ---------------- 7500
Total 7500 gm
2.5 x + 3.25 ( 7500 - x ) = 22500
2.5 x + 24375 - 3.25 x = 22500
2.5 x - 3.25 x = 22500 - -24375
-0.75 x = -1875
/ -0.75
x = 2500 gm P2.50 Tea type I
5000 gm P3.25 Tea type II

m.ananth@hotmail.ca